# Optimizing Multi-modality Trackers via Sensitivity-regularized Tuning

### v2

@@ III-B1 Safeguarding Generalization via Prior Sensitivity
where $D_{u}=D_{0}\bigcup D_{t}$, $\beta>0$ balances the contributions of pre-trained task, and $\mathcal{L}\left(\theta\mid D_{0}\right)$ serves as a regularization term to refine the vanilla objective. Under this joint objective, the pre-trained weights $\theta_{0}$ minimize $\mathcal{L}\left(\theta\mid D_{0}\right)$ and can serve as a local optimum of $\mathcal{L}\left(\theta\mid D_{u}\right)$. When $\theta$ is in the vicinity of $\theta_{0}$, the pre-training loss can be locally approximated by a second-order Taylor expansion:
@@
where $\nabla\mathcal{L}\left(\theta_{0}\right)\approx\mathbf{0}$, and $\mathcal{F}^{\left(\theta_{0}\right)}\in\mathbb{R}^{|\theta|\times|\theta|}$ is expectation of Hessian matrix $\frac{\partial^{2}\mathcal{L}}{\partial\theta_{i}\theta_{j}}$ over dataset $D_{0}$, also called as the Fisher Information Matrix (FIM) [1] at $\theta_{0}$. Principally, it captures the pre-trained tangent space, along which the impact on the pre-trained loss is minimized.
@@
This generalization gap is measured as the weight distance within the Riemannian manifold defined by FIM. Through geometric perspective, FIM yields insights that large deviations from pre-trained tangent space entail a higher risk of generalization degradation [30, 50]. Therefore, FIM serves as a natural and principled prior sensitivity to reflect the erosion of pre-trained knowledge.
Figure 4: Operation-wise prior sensitivity (i.e., eigen-decomposition FIM map) of the pre-trained OSTrack. The high prior sensitivity indicates a tendency to deviate from the pre-trained tangent space, reflecting the disruption of pretrained knowledge.
@@
Empirical studies [14, 37] have shown that the FIM of deep networks is spectrally concentrated, featuring a few large eigenvalues and a long, flat tail, which motivates a low-rank approximation using its leading eigenpairs. Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition, with unit eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$ and eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ sorted $\lambda^{j}_{1}\geq\cdots\geq\lambda^{j}_{K}\geq 0$.
We obtain these leading eigenpairs via Rayleigh-quotient probing [26]. For any non-zero direction $\epsilon$, the Rayleigh quotient is defined as:
@@
Proposition 1 (Eigen-based Approximation Error Bound)
Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition of group $j$, with top-$K$ eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ and eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$. We construct the following approximation:
@@
Proof 1
See Appendix I-A.
Prior Sensitivity Measure. Specifically, we estimate large eigenvalues of FIM using symmetric finite-difference probes [23]. Around $\theta_{0}$ on $D_{0}$, small fixed-radius perturbations expose the directions that maximize the normalized generalization gap, which is equivalent to the Rayleigh quotient of FIM. This procedure systematically maps the pre-trained loss landscape to recover the sharpest generalization shifts and identify the most sensitive parameters. For each group $\theta^{j}$, by sampling various unit directions $\epsilon^{j}\in\mathbb{R}^{|\theta^{j}|}$ from an isotropic Gaussian distribution $\epsilon^{j}\sim\mathcal{N}\left(\mathbf{0},I\right)$ on the weight sphere, we obtain the empirical measures:
@@
where $\mathcal{L}(\theta_{0}^{j})\triangleq\mathcal{L}(\theta_{0}^{j}|D_{0})$, $\rho=1\times 10^{-5}$ is a noise radius. By repeating this search, we obtain an approximation of the top-$K$ eigenvalues. This measure provides a principled yet tractable perspective to assess how each group of the model affects generalization. By considering the eigenvalue magnitudes from different perturbation directions, we condense the prior sensitivity of the $j^{\text{th}}$ operation into a single scalar by distilling its dominant values:
@@
The prior sensitivity is equivalent to FIM in Eq. (6), as presented in Fig. 4. We set $K=10$ in our experiments. In practice, we observe that the leading eigenvalue alone (e.g., $K=1$) provides a reliable indicator of prior sensitivity. Further qualitative results are reported in Section IV-C.
III-B2 Stabilizing Adaptation via Transfer Sensitivity
@@
Characterizing Transfer-Gradient Sparsity. While prior sensitivity addresses forgetting of pre-trained domain, inadaptability in the downstream domain is notably pronounced. Multi-modal tracking often involves diverse and heterogeneous samples exhibiting substantial gaps, such as auxiliary inputs or degraded images. These shifts manifest as highly sparse gradients during fine-tuning process on $D_{t}$, as illustrated in Fig. 5. High sparsity indicates that only a few gradients dominate the updates, while the majority remain nearly static. This phenomenon sheds light on a key factor underlying the limited adaptability of multi-modal trackers and potentially results in temporal oscillatory (see FFT test branch in Fig. 1). In this context, we investigate the elevated adaptation risk induced by sparse gradients and propose a transfer sensitivity aimed at mitigating its impact.
Formally, we quantify the sparsity of the gradient by comparing its $L_{1}$ and $L_{2}$ norms. Let $\mathcal{G}\doteq\nabla_{\theta}\mathcal{L}\left(\theta\mid M\right)$ be the gradient on the current batch $M\subset D_{t}$. We define the sparsity measure: $\rho=\frac{\left\|\mathcal{G}\right\|_{1}}{\sqrt{|\theta|}\left\|\mathcal{G}\right\|_{2}}$. This definition leverages the fact that $L_{1}$ and $L_{2}$ norms characterize different aspects of the vector: while $L_{1}$ reflects the total variation, $L_{2}$ emphasizes concentration [20]. Thus, their ratio serves as an effective proxy for sparsity structure, where a smaller value means higher sparsity. Moreover, noting that single-step gradient magnitudes are typically bounded [35], increasing sparsity (i.e., $\rho:1\rightarrow\frac{1}{\sqrt{|\theta|}}$) can amplify the $L_{2}$ norm:
@@
Figure 5: Instantaneous transfer sensitivity on the VisEvent dataset observed during fine-tuning. The high sparsity in the sensitivity map indicates that only a few gradients (light areas) dominate the updates.
Sparsity-Induced Adaptation Risk. We characterize the tuning volatility as the degree to which performance responds to small parameter perturbations during fine-tuning. Accordingly, we define the following sensitivity function:
@@
By incorporating the linear approximation $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\delta$ into Eq. (11), it can be shown that this risk is proportional to the $L_{2}$ norm of gradient:
@@
Derivations details of Eq. (9)-(12) can be found in Appendix I-B. Thus, we can deduce that sparse gradients exacerbate tuning instability and adaptation risk. This highlights the need to avoid concentrating updates on specific parameters and emphasizes the importance of a balanced transfer gradient.
Transfer Sensitivity Estimation. To formalize the estimation of transfer sensitivity, we adapt the formulation of Eq. (10) to the standard tuning process. In this case, we adjust the parameter $\theta$ via the gradient $\mathcal{G}$ rather than noise perturbations. Therefore, we can derive: $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\mathcal{G}$. Here, the sensitivity function is used to assess the overall adjustment of all parameters during transfer. To further investigate and regularize gradients, we extend the sensitivity to a parameter-wise formulation:
@@
where $s_{n}^{t}$ denotes the transfer sensitivity of the $n^{\text{th}}$ parameter. This formulation allows us to examine the granular impact of parameter changes during the transfer process and rebalance the gradient accordingly.
@@
Step1: Prior Sensitivity Estimation:
@@
 Calculate prior sensitivity $s_{j}^{p}$ via Eq. (8);
Step2: Transfer Sensitivity Estimation and Regularized Tuning:
@@
  Update transfer sensitivity by $s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M_{i}\right)}{\partial\theta_{n}})^{2}$;
  Normalize and combine sensitivity $s_{n}$ by Eq. (14);
@@
Algorithm 1 Sensitivity-Regularized Fine-Tuning
III-C Sensitivity-regularized Tuning
We regularize the learning process by previously discussed sensitivity metrics to derive enhanced trackers. To align with the unified regularization tuning, we rank both the operation-wise prior sensitivity and parameter-wise transfer sensitivity, and then normalize them within the continuous range of $[0,1]$. To harmonize these two sensitivities, we devise a dynamic linear schedule to adjust their weighted combination. At the beginning of training, prior sensitivity plays a dominant role, contributing a weight of $\kappa$ (where $\kappa\in[0,1]$). As training progresses, the influence of transfer sensitivity gradually increases, reaching the same weight $\kappa$ by the end. This ensures that the model initially focuses on retaining pre-trained knowledge and progressively shifts to emphasize training stability. Formally, at each training step $i$, the combined sensitivity is updated as follows:
@@ III-C Sensitivity-regularized Tuning
where $\theta_{n}\in\theta^{j}$, $T$ is the total number of training steps. Finally, we normalize this joint sensitivity within the range of $s_{n}\in[0,0.99]$. To this end, during the training process, parameters that are excessively updated will be penalized based on their sensitivity:
$\theta_{n}^{(i+1)}=\theta_{n}^{(i)}-(1-s_{n})\alpha\frac{\partial\mathcal{L}}{\partial\theta_{n}},\text{s.t. }\theta^{(0)}=\theta_{0},$ (15)
where $\alpha$ is the learning rate. This formulation suggests that more sensitive parameters should retain their previous states to a greater extent, to avoid oscillations or over-adjustments.
The detailed regularization tuning process is given in Algorithm 1. We set the sensitivity harmony coefficient $\kappa=0.6$ in our experiments. Further comprehensive analysis can be found in Table VI.
III-D Learning Objectives
@@
where $\mathcal{L}_{cls}$ is the weighted focal loss for classification, $L_{1}$ loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{iou}$ are employed for bounding box regression, $\lambda_{iou}=2$ and $\lambda_{l_{1}}=5$ are the regularization factors, and all the corresponding settings are the same as [58].
@@
IV Experiments
IV-A Experiment Settings
IV-A1 Benchmark Datasets
To verify the effectiveness and generalization of the proposed method, we conduct comprehensive experiments on seven multi-modal benchmark datasets. For RGB-Event tracking, we use FE108 [61] (76 train / 32 test sequences) captured under degraded conditions, VisEvent [45](500 train / 320 test sequences) covering dynamic outdoor scenes, and CoeSot [42] with 578K image-event pairs (824 train / 528 test sequences). For RGB-Depth tracking, we adopt DepthTrack [56] (152 train / 50 test videos) and VOT-RGBD2022 [22] (127 test sequences). For RGB-Thermal tracking, we use LasHeR [25] (979 train / 245 test sequences) and RGBT234 [24] (234 test videos). All training and evaluation strictly follow established protocols [70, 51]. During fine-tuning, we use the training sets of FE108, VisEvent, and CoeSot for RGB-Event tasks; DepthTrack for RGB-Depth; and LasHeR for RGB-Thermal. The corresponding test sets, along with VOT-RGBD2022 and RGBT234, are used for evaluation.
IV-A2 Evaluation Metrics
Adhering to recognized standards [17, 51, 16, 41], we evaluated tracking performance with the following metrics. For all RGB-Event benchmarks (FE108, VisEvent and CoeSot), We utilize two widely used metrics, i.e., success rate (SR), precision rate (PR). For DepthTrack, precision (Pr) and recall (Re) are used, with F-score ($F=\frac{2Re\cdot Pr}{Re+Pr}$) as the primary measure. For VOT-RGBD2022, we assess accuracy (Acc), robustness (Rob), and expected average overlap (EAO). For LasHeR, success rate (SR) and precision rate (PR) are employed. For RGBT234, maximum success rate (MSR) and maximum precision rate (MPR) served as evaluation metrics. Notably, our SR/PR precisely aligns with SR/PR in ViPT and OneTrack, Suc/Pre in SDSTrack, Success/Precision in UnTrack, or AUC/P in SUTrack. For all metrics, the larger, the better.
IV-A3 Pre-trained Models and Compared Methods
In this paper, we choose three prototypical one-stream RGB-based trackers as pre-trained baselines: OSTrack [58], the most widely adopted model; DropTrack [49] and SUTrack [9], which offer improved generalization. Notably, OSTrack and DropTrack are built upon ViT-B/16 [12], while SUTrack adopts the HiViT-B/16 [64] as the backbone. Following their pre-trained configurations, these models differ in input resolution: OSTrack ($\times128$$\times256$), DropTrack, and SUTrack ($\times192$$\times384$) for template and search regions, respectively. To comprehensively validate the effectiveness of our method, we conduct the following experiments in Table I and Table II. First, we construct strong RGB-based single-modal baselines under full fine-tuning. Then, we evaluate our method under both full and parameter-efficient fine-tuning paradigms. For fair comparison, all methods are grouped according to the pre-trained trackers. Notably, OneTracker [16] and MamTrack [41] adopt non-standard experimental settings and lack open-source implementations. OneTracker uses a $\times384$ pre-trained OSTrack, incompatible with standard benchmarks, so we compare it separately in Table III. MamTrack, with unclear pretraining details but sharing the HiViT-B/16 backbone with SUTrack, is included in SUTrack-based comparisons.
IV-A4 Training Details
We follow the data processing pipeline of SDSTrack [17] across all datasets, converting event data into color-polar event images, with no preprocessing for other auxiliary modalities. The models are trained on 8 NVIDIA 3090Ti GPUs with a batch size of 192 and 30 epochs. Each epoch involves sampling 80k samples. We utilize the AdamW optimizer with a learning rate set to $1\times 10^{-4}$ and a weight decay set to $10^{-4}$.
TABLE I: Quantitative comparison on the RGB-Event datasets. The best are marked with “bold”, and the second best are marked with “underline”.

### v3

@@ 3.2.1 Safeguarding Generalization via Prior Significance
where $D_{u}=D_{0}\bigcup D_{t}$, $\beta>0$ balances the contributions of pre-trained task, and $\mathcal{L}\left(\theta\mid D_{0}\right)$ serves as a regularization term to refine the vanilla objective $\mathcal{L}\left(\theta\mid D_{t}\right)$. Under this joint objective, the pre-trained weights $\theta_{0}$ minimize $\mathcal{L}\left(\theta\mid D_{0}\right)$ and can serve as a local optimum of $\mathcal{L}\left(\theta\mid D_{u}\right)$. When $\theta$ is in the vicinity of $\theta_{0}$, the pre-training loss can be locally approximated by a second-order Taylor expansion:
@@
where $\nabla\mathcal{L}\left(\theta_{0}\right)\approx\mathbf{0}$, and $\mathcal{F}^{\left(\theta_{0}\right)}\in\mathbb{R}^{|\theta|\times|\theta|}$ is expectation of Hessian matrix $\frac{\partial^{2}\mathcal{L}}{\partial\theta_{i}\theta_{j}}$ over dataset $D_{0}$, also called as the Fisher Information Matrix (FIM) Amari et al. (2019) at $\theta_{0}$. Principally, it captures the pre-trained tangent space, along which the impact on the pre-trained loss is minimized.
@@
This generalization gap is measured as the weight distance within the Riemannian manifold defined by FIM. Through geometric perspective, FIM yields insights that large deviations from pre-trained tangent space entail a higher risk of generalization degradation (Liu et al., 2022; Wu et al., 2024a). Therefore, FIM serves as a natural and principled prior significance to reflect the erosion of pre-trained knowledge.
Figure 4: Operation-wise prior parameter significance (i.e., eigen-decomposition of the FIM) for the pre-trained OSTrack on the source datasets. The high prior significance indicates a tendency to deviate from the pre-trained tangent space, reflecting the disruption of pretrained knowledge.
@@
Empirical studies (Ghorbani et al., 2019; Rame et al., 2022) have shown that the FIM of deep networks is spectrally concentrated, featuring a few large eigenvalues and a long, flat tail, which motivates a low-rank approximation using its leading eigenpairs. Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition, with unit eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$ and eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ sorted $\lambda^{j}_{1}\geq\cdots\geq\lambda^{j}_{K}\geq 0$.
We obtain these leading eigenpairs via Rayleigh-quotient probing (Li, 2015). For any non-zero direction $\epsilon$, the Rayleigh quotient is defined as:
@@
Proposition 1(Eigen-based Approximation Error Bound.)
Let $\tilde{\mathcal{F}}^{(\theta_{0}^{j})}={(V^{j})}\Lambda^{j}{(V^{j}})^{T}$ be the eigen-decomposition of parameter group $j$, with top-$K$ eigenvalues $\Lambda^{j}=\mathrm{diag}(\lambda^{j}_{1},\dots,\lambda^{j}_{K})$ and eigenvectors $V^{j}=[v^{j}_{1},\dots,v^{j}_{K}]$. We construct the following approximation:
@@
Proof
See Appendix B.
Prior Significance Measurement. Specifically, we estimate large eigenvalues of FIM using symmetric finite-difference probes (LeVeque, 1998). Around $\theta_{0}$ on $D_{0}$, small fixed-radius perturbations expose the directions that maximize the normalized generalization gap, which is equivalent to the Rayleigh quotient of FIM. This procedure systematically maps the pre-trained loss landscape to recover the sharpest generalization shifts and identify the most significant parameters. For each group $\theta^{j}$, by sampling various unit directions $\epsilon^{j}\in\mathbb{R}^{|\theta^{j}|}$ from an isotropic Gaussian distribution $\epsilon^{j}\sim\mathcal{N}\left(\mathbf{0},I\right)$ on the weight sphere, we obtain the empirical measures:
@@
where $\mathcal{L}(\theta_{0}^{j})\triangleq\mathcal{L}(\theta_{0}^{j}|D_{0})$, $\rho=1\times 10^{-5}$ is a noise radius. By repeating this search, we obtain an approximation of the top-$K$ eigenvalues. This measure provides a principled yet tractable perspective to assess how each group of the model affects generalization. By considering the eigenvalue magnitudes from different perturbation directions, we condense the prior significance of the $j^{\text{th}}$ operation into a single scalar by distilling its dominant values:
@@
The prior significance is equivalent to FIM in Eq. (6), as presented in Fig. 4. We set $K=10$ in our experiments. In practice, we observe that the leading eigenvalue alone (e.g., $K=1$) provides a reliable indicator of prior significance. Further qualitative results are reported in Section 4.3.
3.2.2 Stabilizing Adaptation via Transfer Significance
@@
Characterizing Transfer-Gradient Sparsity. While prior significance addresses forgetting of pre-trained domain, inadaptability in the downstream domain remains notably pronounced. Specifically, multi-modal tracking often involves diverse and heterogeneous samples exhibiting substantial gaps, such as auxiliary inputs or degraded images. These shifts manifest as highly sparse gradients during fine-tuning process on $D_{t}$, as illustrated in Fig. 5. High sparsity indicates that only a few gradients dominate the updates, while the majority remain nearly static. This phenomenon sheds light on a key factor underlying the limited adaptability of multi-modal trackers and potentially results in temporal oscillatory (see FFT test branch in Fig. 1). In this context, we investigate the elevated adaptation risk induced by sparse gradients and propose a transfer significance aimed at mitigating its impact.
Formally, we quantify the sparsity of the gradient by comparing its $L_{1}$ and $L_{2}$ norms. Let $\mathcal{G}\doteq\nabla_{\theta}\mathcal{L}\left(\theta\mid M\right)$ be the gradient on the current batch $M\subset D_{t}$. We define the sparsity measure: $\rho=\frac{\left\|\mathcal{G}\right\|_{1}}{\sqrt{|\theta|}\left\|\mathcal{G}\right\|_{2}}$. This definition leverages the fact that $L_{1}$ and $L_{2}$ norms characterize different aspects of the gradient vector: while $L_{1}$ reflects the total variation, $L_{2}$ emphasizes concentration Hurley and Rickard (2009). Thus, their ratio serves as an effective proxy for sparsity structure, where a smaller value means higher sparsity. Moreover, noting that single-step gradient magnitudes are typically bounded Nesterov (2013), increasing sparsity (i.e., $\rho:1\rightarrow\frac{1}{\sqrt{|\theta|}}$) can amplify the $L_{2}$ norm of the gradient vector:
@@
Figure 5: Instantaneous transfer parameter significance on the VisEvent dataset observed during fine-tuning. The high sparsity in the significance map indicates that only a few gradients (light areas) dominate the updates.
Sparsity-Induced Adaptation Risk. We characterize the tuning volatility as the degree to which performance responds to small parameter perturbations during fine-tuning. Accordingly, we define the following response function:
@@
By incorporating the linear approximation $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\delta$ into Eq. (11), it can be shown that this adaptation risk scales with the $L_{2}$ norm of the gradient vector:
@@
Derivations details of Eqs. (9)-(12) are provided in Appendix B. Thus, we can deduce that sparse gradients exacerbate tuning instability and adaptation risk. This highlights the need to avoid concentrating updates on specific parameters and emphasizes the importance of a balanced transfer gradient.
Transfer Significance Estimation. To formalize the estimation of transfer significance, we adapt the formulation of Eq. (10) to the standard fine-tuning process. In this case, we adjust the parameter $\theta$ via the gradient $\mathcal{G}$ rather than noise perturbations. Therefore, we can derive: $S\left(\theta,\delta\mid M\right)\approx\mathcal{G}^{T}\mathcal{G}$. Here, the significance function is used to assess the overall adjustment of all parameters during transfer. To further investigate and regularize gradients, we extend the transfer significance to a parameter-wise formulation:
@@
where $s_{n}^{t}$ denotes the transfer significance of the $n^{\text{th}}$ parameter. This formulation allows us to examine the granular impact of parameter changes during the transfer process and rebalance the gradient accordingly.
Algorithm 1 Significance-Regularized Fine-Tuning
@@
Step1: Prior Significance Estimation:
@@
 Calculate prior significance $s_{j}^{p}$ via Eq. (8);
Step2: Transfer Significance Estimation and Regularized Tuning:
@@
  Update transfer significance by $s_{n}^{t}=(\frac{\partial\mathcal{L}\left(\theta\mid M_{i}\right)}{\partial\theta_{n}})^{2}$;
  Normalize and combine significance $s_{n}$ by Eq. (14);
@@
3.3 Significance-Regularized Fine-Tuning
We regularize the learning process by previously discussed significance metrics to derive enhanced multi-modality trackers. To align with the unified regularization tuning, we rank both the operation-wise prior significance and parameter-wise transfer significance, and then normalize them within the continuous range of $[0,1]$. To harmonize these two parameter significance, we devise a dynamic linear schedule to adjust their weighted combination. At the beginning of training, prior significance plays a dominant role, contributing a weight of $\kappa$ (where $\kappa\in[0,1]$). As training progresses, the influence of transfer significance gradually increases, reaching the same weight $\kappa$ by the end. This ensures that the model initially focuses on retaining pre-trained knowledge and progressively shifts to emphasize training stability. Formally, at each training step $i$, the combined parameter significance is updated as follows:
@@ 3.3 Significance-Regularized Fine-Tuning
where $\theta_{n}\in\theta^{j}$, $T$ is the total number of training steps. Finally, we normalize this joint significance within the range of $s_{n}\in[0,0.99]$. To this end, during the training process, parameters that are excessively updated will be penalized based on their significance:
$\theta_{n}^{(i+1)}=\theta_{n}^{(i)}-(1-s_{n})\alpha\frac{\partial\mathcal{L}}{\partial\theta_{n}},\quad{s.t.}\quad\theta^{(0)}=\theta_{0},$ (15)
where $\alpha$ is the learning rate. This formulation suggests that more significant/sensitive parameters should retain their previous states to a greater extent, to avoid oscillations or over-adjustments.
The detailed regularization tuning process is given in Algorithm 1. We set the significance harmony coefficient $\kappa=0.6$ in our experiments. Further comprehensive analysis can be found in Table 6.
Discussion. Unlike prior works that rely on static “sensitivity” to rank parameters for sparse tuning, SRFT introduces a fundamentally distinct notion of “significance” — a hybrid, dynamic estimate combining pre-trained Fisher eigen-structure and instantaneous gradient sparsity. Instead of selecting the most sensitive parameters for exclusive updating, SRFT prioritizes penalizing them to suppress excessive updates, without structural constraints or discontinuous optimization. This regularization-centric design yields a smooth, adaptive update path tailored for cross-modal transfer, diverging from existing sparsity- or retention-oriented approaches. Additional comparative experiments can be found in Appendix C.
3.4 Learning Objectives
@@
where $\mathcal{L}_{cls}$ is the weighted focal loss for classification, $L_{1}$ loss $\mathcal{L}_{1}$ and GIoU loss $\mathcal{L}_{iou}$ are employed for bounding box regression, $\lambda_{iou}=2$ and $\lambda_{l_{1}}=5$ are the regularization factors, and all the corresponding settings are the same as (Ye et al., 2022).
@@
4 Experiments
4.1 Experiment Settings
4.1.1 Benchmark Datasets
To verify the effectiveness and generalization of the proposed method, we conduct comprehensive experiments on seven multi-modal benchmark datasets. For RGB-Event tracking, we use FE108 (Zhang et al., 2021a) (76 train / 32 test sequences) captured under degraded conditions, VisEvent (Wang et al., 2023)(500 train / 320 test sequences) covering dynamic outdoor scenes, and CoeSot (Tang et al., 2022) with 578K image-event pairs (824 train / 528 test sequences). For RGB-Depth tracking, we adopt DepthTrack (Yan et al., 2021b) (152 train / 50 test videos) and VOT-RGBD2022 (Kristan et al., 2022) (127 test sequences). For RGB-Thermal tracking, we use LasHeR (Li et al., 2020) (979 train / 245 test sequences) and RGBT234 (Li et al., 2019) (234 test videos). All training and evaluation strictly follow established protocols (Zhu et al., 2023a; Wu et al., 2024b). During fine-tuning, we use the training sets of FE108, VisEvent, and CoeSot for RGB-Event tasks; DepthTrack for RGB-Depth tasks; and LasHeR for RGB-Thermal tasks. The corresponding test sets, along with VOT-RGBD2022 and RGBT234, are used for evaluation.
Table 1: Quantitative comparison on the RGB-Event datasets. The best are marked with “bold”, and the second best are marked with “underline”.